OpenStudy (anonymous):

Team Bear’s kite is 4.5 ft long and 3.5 ft across. How much fabric does it take to make their kite? A. 15.75 B. 8 C. 16 D. 7.875

OpenStudy (anonymous):

The surface are of the kite = (1/2)*(4.5)*(3.5) = 7.875 square feet

OpenStudy (anonymous):

area of the kite=1/2*d1*d2

OpenStudy (crashonce):

what @dg2 means is that the area is 1/2 * multiplication of the diagonals

OpenStudy (anonymous):

yes